OpenStudy (gfdgdgdgfhdf):

Solve for the variable: 11p + 15 = 81

OpenStudy (madisonrenee):

6

OpenStudy (madisonrenee):

81-15= 66, 66/11=6

OpenStudy (gfdgdgdgfhdf):

ty

OpenStudy (stbest):

^^^ 111 6*11=66+15=81